### v1

# Multigroup Radiation Diffusion on a Moving Mesh: Implementation in rich and Application to Tidal Disruption Events

Abstract
Radiation-hydrodynamics (RHD) determines the bulk evolution and observable emission in a wide variety of high-energy astrophysical phenomena. Due to their complexity, RHD problems must usually be studied through numerical simulation. We have extended the publicly available rich code, which previously solved the equations of RHD in the limit of grey flux-limited diffusion (FLD), to operate with a multigroup FLD solver. rich is a semi-Lagrangian code that solves the equations of RHD on an unstructured moving mesh, and is the first multigroup RHD moving mesh code, making it uniquely applicable to problems with extreme dynamic range and dynamically important radiation forces. We validate our multigroup module against multiple analytic benchmarks, including a novel test of the RHD Doppler term. The computational efficiency of the code is aided by a novel scheme to accelerate convergence in optically thick cells. Finally, we apply multigroup rich in a pilot study of a stellar tidal disruption event (TDE), using a $10^{4}M_{\odot}$ intermediate-mass black hole. Our simulations self-consistently produce a bright early-time X-ray flash prior to peak optical/UV light, in qualitative agreement with post-processing of (grey) rich simulations of supermassive black hole TDEs, as well as X-ray observations of the TDE AT 2022dsb.
I Introduction
Radiation-hydrodynamics (RHD) is central to a multitude of astrophysical processes. Energy transport in stellar interiors (Kippenhahn et al., 2013), and the flow of matter in (Yuan and Narayan, 2014; Jiang et al., 2019) and out (Fabian, 1999; Thompson et al., 2015) of active galactic nuclei exemplify quasi-stationary problems in RHD, for which it is possible to construct approximate analytic or semi-analytic solutions. Conversely, other important astrophysical phenomena occur in the dynamical limit of RHD, such as the cooling and fragmentation of gas in star-forming regions (Krumholz et al., 2007b; Grudić et al., 2021), or the evolution of complex transients around black holes (Bonnerot et al., 2021; Steinberg and Stone, 2024; Brutman et al., 2024; Huang et al., 2025a). For dynamical problems such as these, it is uncommon for analytic methods to be practical, and the natural theoretical approach is to discretize and simulate the equations of RHD on a computer.
A variety of numerical methods exist to simulate problems in radiation hydrodynamics (Castor, 2004). Standard hydrodynamics solvers can be coupled to various discretizations of the equation of radiation transfer. The most common approach, flux-limited diffusion (FLD), evolves a one-moment closure of the transfer equation, assuming that radiation transfer is diffusive in optically thick regimes, while using a flux limiter to approximate free streaming motion in optically thin regions (Levermore and Pomraning, 1981). FLD is generally an efficient approach to RHD; other algorithms, such as the M1 two-moment closure (Levermore, 1984; Aubert and Teyssier, 2008; Sądowski et al., 2013), the variable Eddington tensor method (Stone et al., 1992; Hayes and Norman, 2003; Jiang et al., 2012), Monte Carlo schemes (Fleck and Cummings (1971); Gentile (2001); Wollaber (2016); Noebauer and Sim (2019); Steinberg and Heizler (2022a, 2023)) or $S_{N}$ and $P_{N}$ deterministic schemes (Brunner (2002); McClarren (2010); Till et al. (2018)), will usually attain greater physical accuracy at the expense of greater computational cost. Aside from the choice of discretization, a second axis of differentiation between RHD solvers is the treatment of photon spectra. Some algorithms work in the grey, or single-frequency, approximation, while other multigroup solvers operate on multiple different bins of photons, which propagate differently due to the frequency-dependent opacity of the matter they interact with.
In this paper, we take the existing rich radiation hydrodynamics code and upgrade its radiation transfer module from grey to multigroup FLD. rich is a publicly available code developed originally for finite-volume semi-Lagrangian (moving-mesh) hydrodynamics (Yalinewich et al., 2015; Steinberg et al., 2015, 2016). Moving mesh algorithms for hydrodynamics add computational expense (e.g. repeated tessellation of an unstructured mesh) but can avoid common pitfalls of finite mass schemes while retaining their strengths in problems with enormous dynamic range or highly supersonic flow. Cold flow accretion into galaxies (Nelson et al., 2013), dynamically cold shearing flows (Zier and Springel, 2022), and stellar tidal disruption events (Steinberg and Stone, 2024) are examples of problems that pose severe computational challenges for both Eulerian, static mesh schemes and finite mass approaches, but which have been fruitfully tackled using moving-mesh codes.
Indeed, tidal disruption events (TDEs) are one of the main use cases for rich (Krolik et al., 2020), both in its original, purely hydrodynamic form, and in its newer (grey) RHD version. rich has already been used to perform notable global simulations of TDEs, including the first end-to-end simulation of an astrophysically realistic disruption (Steinberg and Stone, 2024), and the most detailed series of global TDE convergence tests to date (Martire et al., 2025). However, these studies have been limited by the grey radiation transfer module in the existing version of rich, which has prevented self-consistent modeling of emergent spectra. In past rich simulations of TDEs, spectral properties have either been ignored, or have been computed through simplified post-processing schemes (Steinberg and Stone, 2024). The primary goal of this paper is to implement multigroup FLD radiation transfer in rich, enabling more sophisticated simulations of high-energy astrophysical processes, including TDEs, stellar collisions (Brutman et al., 2024) and novae (Steinberg and Metzger, 2018). A secondary goal of this paper is to properly document the rich FLD algorithm, which had not previously been written up in a code paper. Only a small number of moving-mesh radiation hydrodynamics codes exist, such as AREPO-RT (Kannan et al., 2019), MANGA (Chang et al., 2020), and AREPO-IDORT (Ma et al., 2025), all of which operate in the grey limit of radiation transfer, making this the first multigroup moving-mesh RHD code.
In Section II, we review the monochromatic, comoving, flux‑limited diffusion (FLD) equation and its coupling to material energy and momentum. Section III derives the multigroup FLD equations from the monochromatic form. Section IV details the rich implementation, including the spatial discretization and the fully implicit time integration for the coupled material–radiation system, and outlines the solution strategy for the resulting large linear system. Section V presents verification tests (Marshak waves, a radiative shock, comparisons with multigroup Monte Carlo, and a Doppler‑term test) and an application to a tidal disruption event (TDE), demonstrating the method’s applicability to astrophysical flows. We summarize in Section VI and discuss future work.
II The Monochromatic Radiation Diffusion Equation
We adopt the frequency‑dependent, comoving‑frame flux‑limited diffusion (FLD) equation with elastic scattering (i.e. scattering contributes to the transport opacity but does not exchange energy with the material). The spectral radiation energy density $E_{\nu}(\mathbf{x},t)$, which has units of energy per unit photon-energy per unit volume, evolves according to (Castor 2004; Mihalas and Weibel-Mihalas 1999; Winslow 1995):
$\begin{split}&\frac{\partial E_{\nu}}{\partial t}+\nabla\cdot\left(\vec{v}E_{\nu}\right)=-\frac{1-R_{2,\nu}}{2}E_{\nu}\nabla\cdot\vec{v}+\nabla\cdot\left(D_{\nu}\nabla E_{\nu}\right)\\ &+c\kappa_{\nu}\left(b_{\nu}U_{m}-E_{\nu}\right)+\frac{\partial\left(\nu E_{\nu}\left(1-R_{2,\nu}\right)/2\right)}{\partial\nu}\nabla\cdot\vec{v},\end{split}$ (II.1)
where $\vec{v}$ is the fluid velocity, $\kappa_{\nu}$ is the monochromatic absorption opacity, $\nu$ is the radiation frequency (in energy units), $U_{m}\equiv aT^{4}$ is the blackbody radiation energy density at the material temperature, $a$ is the radiation constant, and $T$ is the material temperature. The normalized Planck function is
$b_{\nu}\equiv B_{\nu}(T)/U_{m}$ (II.2)
such that $\int_{0}^{\infty}b_{\nu}\,d\nu=1$, and
$B_{\nu}(T)=\frac{8\pi\nu^{3}}{h^{3}c^{3}}\frac{1}{\exp\left(\frac{\nu}{k_{B}T}\right)-1}$ (II.3)
is the Planck spectrum in energy-density form. We note that $B_{\nu}$ here has units of energy density per photon-energy, and is normalized such that $\intop_{0}^{\infty}B_{\nu}(T)d\nu=U_{m}$. In the flux–limited diffusion approximation, the effective radiative diffusion coefficient $D_{\nu}$ is given by
$D_{\nu}=\frac{c\lambda_{\nu}}{\kappa_{t,\nu}},$ (II.4)
where $\kappa_{t,\nu}$ is the total opacity (absorption + scattering) of a photon of frequency $\nu$, and $\lambda_{\nu}$ is a flux limiter. We employ the Levermore–Pomraning flux-limiter (Levermore and Pomraning, 1981), given by:
$\begin{split}&\lambda_{\nu}=\frac{1}{R_{\nu}}\left(\coth{R_{\nu}}-\frac{1}{R_{\nu}}\right)\\ &R_{\nu}=\frac{\left|\nabla E_{\nu}\right|}{\kappa_{t,\nu}E_{\nu}},\ R_{2,\nu}=\lambda_{\nu}+\lambda_{\nu}^{2}R_{\nu}^{2}\end{split}$ (II.5)
The terms in Eq. II.1 encode different forms of work and energy transport, and their physical meanings can be understood in the following ways: $\nabla\cdot\left(\vec{v}E_{\nu}\right)$ describes comoving advection of the radiation field in the fluid flow, $\frac{1}{2}(1-R_{2,\nu})E_{\nu}\nabla\cdot\vec{v}$ describes a mechanical $P{\rm d}V$ exchange, $\nabla\cdot\left(D_{\nu}\nabla E_{\nu}\right)$ describes diffusion of radiation energy, $c\kappa_{\nu}\left(b_{\nu}U_{m}-E_{\nu}\right)$ couples the radiation energy density to the material energy density through emission/absorption, and $\frac{\partial}{\partial\nu}\left(\nu E_{\nu}\left(1-R_{2,\nu}\right)/2\right)\nabla\cdot\vec{v}$ encodes the Doppler shift.
The total material energy density $E_{m}$ couples to the radiation field via the terms (Castor 2004; Mihalas and Weibel-Mihalas 1999):
$\begin{split}&\frac{\partial E_{m}}{\partial t}+\nabla\cdot\left[\left(E_{m}+P\right)\vec{v}\right]=\\ &-\vec{v}\cdot\intop_{0}^{\infty}\lambda_{\nu}\nabla E_{\nu}d\nu-c\intop_{0}^{\infty}\kappa_{\nu}(b_{\nu}U_{m}-E_{\nu})d\nu.\end{split}$ (II.6)
The radiation force enters the momentum density equation via the term
$\frac{\partial(\rho\vec{v})}{\partial t}+\nabla\cdot\left(\rho\vec{v}\vec{v}\right)=-\nabla P-\intop_{0}^{\infty}\lambda_{\nu}\nabla E_{\nu}d\nu.$ (II.7)
rich advances Eqs. (II.1)–(II.7) with operator splitting, and the purely hydrodynamic terms $\nabla\cdot[(E_{m}+P)\vec{v}]$, $\nabla\!\cdot(\rho\vec{v}\vec{v})$, and $\nabla P$ are updated in the hydro step and are unaffected by the distinction between grey and multigroup FLD. In what follows we focus on the radiation–matter system, omitting the hydro-only terms from the notation.
III The Multigroup Diffusion Equation
We partition the spectrum into a grid of $G$ groups with interfaces at $\nu_{1/2}<\cdots<\nu_{G+1/2}$. The group centers introduce a degree of freedom, and we define them using the arithmetic mean $\nu_{g}\equiv(\nu_{g-1/2}+\nu_{g+1/2})/2$.
We adopt the group-integrated radiation variables $E_{g}$ and $b_{g}$ between interfaces $\nu_{g\pm 1/2}$, centered at $\nu_{g}$:
$E_{g}\;\equiv\;\int_{\nu_{g-1/2}}^{\nu_{g+1/2}}E_{\nu}\,d\nu,\qquad b_{g}\;\equiv\;\int_{\nu_{g-1/2}}^{\nu_{g+1/2}}b_{\nu}\,d\nu,$ (III.1)
The group Planck integrals $b_{g}$ are approximated using a truncated series (see appendix A). We assume that the frequency grid covers the Planckian well for all relevant material temperatures, that is
$\sum_{g=1}^{G}b_{g}(T)\approx 1$
Integrating Eq. II.1 over the range $\nu_{g-1/2},\nu_{g+1/2}]$ yields the multigroup radiation diffusion equation:
$\begin{split}&\frac{\partial E_{g}}{\partial t}+\nabla\cdot\left(\vec{v}E_{g}\right)=-\frac{1-R_{2,g}}{2}E_{g}\nabla\cdot\vec{v}\\ &+\nabla\cdot\left(D_{g}\nabla E_{g}\right)+c\kappa_{P,g}\left(b_{g}U_{m}-E_{g}\right)\\ &+\nabla\cdot\vec{v}\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\frac{\partial\left(\nu E_{\nu}(1-R_{2,\nu})/2\right)}{\partial\nu}d\nu,\end{split}$ (III.2)
where the diffusion coefficient in group $g$ is:
$D_{g}\equiv\frac{c\lambda_{g}}{\kappa_{R,g}},$ (III.3)
with the group flux limiter:
$\begin{split}&\ \lambda_{g}=\frac{1}{R_{g}}\left(\coth{R_{g}}-\frac{1}{R_{g}}\right)\\ &R_{g}=\frac{\left|\nabla E_{g}\right|}{\kappa_{R,g}E_{g}},\ R_{2,g}=\lambda_{g}+\lambda_{g}^{2}R_{g}^{2},\end{split}$ (III.4)
and the group average Rosseland mean opacity:
$\kappa_{R,g}^{-1}=\frac{\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\kappa_{t,\nu}^{-1}\left[\partial B_{\nu}/\partial T\right]d\nu}{\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\partial B_{\nu}/\partial Td\nu}.$ (III.5)
The average group absorption ("Planck") opacity is:
$\kappa_{P,g}=\frac{\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\kappa_{\nu}B_{\nu}(T)d\nu}{\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}B_{\nu}(T)d\nu}.$ (III.6)
Splitting the terms in Eq. II.6 over the energy groups gives
$\frac{\partial E_{m}}{\partial t}=-\vec{v}\cdot\sum_{g^{\prime}}\lambda_{g^{\prime}}\nabla E_{g^{\prime}}+c\left(\sum_{g^{\prime}}\kappa_{P,g^{\prime}}E_{g^{\prime}}-\kappa_{P}U_{m}\right)$ (III.7)
where
$\kappa_{P}=\sum_{g^{\prime}}\kappa_{P,g^{\prime}}b_{g^{\prime}},$ (III.8)
is the total Planck opacity. Here and below we use a primed index $g^{\prime}$ for group sums to distinguish the dummy summation index from the specific group $g$ appearing in the multigroup transport equation, this avoids ambiguity later when substituting the material–radiation coupling back into Eq. (III.2). We separate the material energy equation into an internal‑energy update,
$\frac{\partial u_{m}}{\partial t}=c\left(\sum_{g^{\prime}}\kappa_{P,g^{\prime}}E_{g^{\prime}}-\kappa_{P}U_{m}\right)$ (III.9)
and a kinetic‑energy exchange:
$\frac{\partial K}{\partial t}=-\vec{v}\cdot\sum_{g^{\prime}}\lambda_{g^{\prime}}\nabla E_{g^{\prime}}.$ (III.10)
Similarly, splitting the momentum coupling (Eq. II.7) over the energy groups yields
$\frac{\partial(\rho\vec{v})}{\partial t}=-\sum_{g^{\prime}}\lambda_{g^{\prime}}\nabla E_{g^{\prime}}.$ (III.11)
Equations III.2-III.11 define the coupled multigroup radiation diffusion equations (modulo the hydro-only terms).
IV Implementation in rich
rich first advances the Euler equations using a second-order Godunov scheme, then moves the Voronoi mesh according to the velocity field from the hydrodynamic step (Yalinewich et al., 2015), and finally solves the radiation–transfer equations, either in the mixed-frame grey diffusion approximation (Steinberg and Stone, 2024) or in the multigroup diffusion approximation.
The multigroup diffusion step is itself split into three substages.
  1. 1.
    Advection: we apply the comoving advection contribution $\nabla\cdot\left(\vec{v}E_{g}\right)$ using the mass fluxes from the solution of the Riemann problem. This is performed during the hydro step.
  2. 2.
    Radiation Step: using an implicit finite-volume scheme, we solve the remaining terms of the multigroup system: mechanical $P\mathrm{d}V$ work, diffusion, emission/absorption, and Doppler shifts. Taken together, the remaining terms define the following differential equation for each group:
    $\begin{split}&\frac{\partial E_{g}}{\partial t}=-\frac{1-R_{2,g}}{2}E_{g}\nabla\cdot\vec{v}\\ &+\nabla\cdot\left(D_{g}\nabla E_{g}\right)+c\kappa_{P,g}\left(b_{g}U_{m}-E_{g}\right)\\ &+\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\frac{\partial\left(\nu E_{\nu}(1-R_{2,\nu})/2\right)}{\partial\nu}d\nu\nabla\cdot\vec{v}.\end{split}$ (IV.1)
    This is solved together with the material energy equation (Eq. III.9).
  3. 3.
    Momentum/Kinetic energy: we update the momentum and kinetic energy of the gas using the new radiation field.
IV.1 Discretization of the Multigroup Equation
Integrating Eq. IV.1 over the volume $V_{i}$ of cell $i$ yields:
$\begin{split}&V_{i}\frac{\partial E_{g,i}}{\partial t}=-\frac{1-R_{2,g,i}}{2}E_{g,i}\intop_{V_{i}}\nabla\cdot\vec{v}dV\\ &+\intop_{V_{i}}\nabla\cdot\left(D_{g}\nabla E_{g}\right)dV+cV_{i}\kappa_{P,g,i}\left(b_{g,i}U_{m,i}-E_{g,i}\right)\\ &+\intop_{\nu_{g-1/2}}^{\nu_{g+1/2}}\frac{\partial\left(\nu E_{\nu,i}(1-R_{2,\nu,i})/2\right)}{\partial\nu}d\nu\intop_{V_{i}}\nabla\cdot\vec{v}dV.\end{split}$ (IV.2)
Here $E_{g,i}=\frac{1}{V_{i}}\intop_{V_{i}}E_{g}dV$ is the volume-averaged group radiation energy density. Using the divergence theorem, we can write
$\intop_{V_{i}}\nabla\cdot\vec{v}dV=\intop_{\partial V_{i}}\vec{v}\cdot d\vec{A}=\sum_{j}^{\text{neighbors}}\vec{A}_{ij}\cdot\vec{v}_{ij},$ (IV.3)
where $\partial V_{i}$ is the boundary of cell $i$, the sum is across all neighbors $j$ of cell $i$, $\vec{v}_{ij}=\frac{\vec{v}_{i}+\vec{v}_{j}}{2}$ is the face-averaged velocity, and $\vec{A}_{ij}$ points in the direction of the outward face normal and has a magnitude equal to the area of the face bordering both cell $i$ and cell $j$.
For the diffusion term,
$\intop_{V_{i}}\nabla\cdot\left(D_{g}\nabla E_{g}\right)dV=\sum_{j}^{\text{neighbors}}\vec{A}_{ij}\cdot\left(D_{g,ij}\nabla E_{g,ij}\right),$ (IV.4)
with the face averaged quantities defined as:
$\begin{split}&T_{ij}=\left(\frac{T_{i}^{4}+T_{j}^{4}}{2}\right)^{1/4},\ D_{g,ij}=\frac{c\lambda_{g,ij}}{\kappa_{R,g,ij}}\\ \ &\kappa_{R,g,ij}=\frac{2}{\kappa_{R,g,i}^{-1}+\kappa_{R,g,j}^{-1}},\ R_{g,ij}=\frac{\left|\nabla E_{g,ij}\right|}{\kappa_{R,g,ij}E_{g,ij}}\\ &E_{g,ij}=\frac{E_{g,i}+E_{g,j}}{2}.\end{split}$ (IV.5)
In the calculation of $\kappa_{R,g,ij}$, $\kappa_{R,g,i}$ and $\kappa_{R,g,j}$, the temperature $T_{ij}$ was used in tandem with their respective cell densities (Till, 2020).
The face-centered energy density gradient is calculated using the center-of-mass position $\overrightarrow{CM}_{i}$:
$\nabla E_{g,ij}=\frac{E_{g,j}-E_{g,i}}{\left\|\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right\|^{2}}\left(\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right).$ (IV.6)
IV.1.1 Doppler Term
In this subsection we detail our treatment of the Doppler term,
$\nabla\cdot\vec{v}\frac{\partial\left(\nu E_{\nu}(1-R_{2,\nu})/2\right)}{\partial\nu}.$ (IV.7)
Defining the negative divergence of the velocity as $\alpha\equiv-\nabla\cdot\vec{v}$ and integrating Eq. IV.7 over group $g$ results in
$\begin{split}&\alpha\nu_{g+1/2}\left(\frac{1-R_{2,\nu_{g+1/2}}}{2}\right)E_{\nu_{g+1/2}}\\ &-\alpha\nu_{g-1/2}\left(\frac{1-R_{2,\nu_{g-1/2}}}{2}\right)E_{\nu_{g-1/2}}\end{split}$ (IV.8)
The interface values (in frequency space) $E_{\nu_{g\pm 1/2}}$ are reconstructed with an upwind, slope–limited scheme. Notice that $E_{\nu_{g\pm 1/2}}$ are energy density per unit frequency. The left slope $m_{L,g}$ and the right slope $m_{R,g}$ are:
$\begin{split}&m_{L,g}=\frac{E_{\nu_{g}}-E_{\nu_{g-1}}}{\Delta\nu_{g-1}}=\frac{{E_{g}}/{\Delta_{g}}-{E_{g-1}}/{\Delta_{g-1}}}{\Delta\nu_{g-1}}\\ &m_{R,g}=\frac{E_{\nu_{g+1}}-E_{\nu_{g}}}{\Delta\nu_{g}}=\frac{{E_{g+1}}/{\Delta_{g+1}}-{E_{g}}/{\Delta_{g}}}{\Delta\nu_{g}}.\end{split}$ (IV.9)
Here $\Delta_{g}=\nu_{g+1/2}-\nu_{g-1/2}$ is the energy group width and $\Delta\nu_{g}=\nu_{g+1}-\nu_{g}$ is the energy difference between the group centers.
The slope-limiter $\phi(r_{g})$ is a function of the group slope ratio $r_{g}=\frac{m_{L,g}}{m_{R,g}}$. In our implementation we use the Superbee slope-limiter (Roe 1986, see section V.3.2), but other limiters can be used as well. The Superbee slope-limiter is given by
$\phi(r)=\max\left(0,\min\left(2r,1\right),\min\left(r,2\right)\right).$ (IV.10)
Since compression ($\alpha\geq 0,\ \nabla\cdot v\leq 0$) blue-shifts the photons (i.e. advects them to higher $\nu$), and expansion red-shifts the photons (i.e. advects to lower $\nu$), the upwind direction switches sides. More concretely: if we consider the $\alpha\geq 0,\ \nabla\cdot v\leq 0$ case, the energy groups get blue-shifted, thus the upwind direction is to the lower energies, and $E_{\nu_{g\pm 1/2}}$ are reconstructed by interpolation from the left side,
$\begin{split}&E_{\nu_{g+1/2}}={E_{g}}/{\Delta_{g}}+\frac{\Delta_{g}\phi\left(r_{g}\right)}{2\Delta\nu_{g}}\left(\frac{E_{g+1}}{\Delta_{g+1}}-\frac{E_{g}}{\Delta_{g}}\right)\\ &E_{\nu_{g-1/2}}=\frac{E_{g-1}}{\Delta_{g-1}}+\frac{\Delta_{g-1}\phi\left(r_{g-1}\right)}{2\Delta\nu_{g-1}}\left(\frac{E_{g}}{\Delta_{g}}-\frac{E_{g-1}}{\Delta_{g-1}}\right).\end{split}$ (IV.11)
Inserting the above into the Doppler term,
$\begin{split}&\alpha\nu_{g+1/2}\left(\frac{1-R_{2,\nu_{g+1/2}}}{2}\right)\Bigg[\frac{E_{g}}{\Delta_{g}}\\ &\qquad\qquad\quad\quad\quad\quad+\frac{\Delta_{g}}{2}\phi\left(r_{g}\right)\frac{{E_{g+1}}/{\Delta_{g+1}}-{E_{g}}/{\Delta_{g}}}{\Delta\nu_{g}}\Bigg]\\ &-\alpha\nu_{g-1/2}\left(\frac{1-R_{2,\nu_{g-1/2}}}{2}\right)\Bigg[\frac{E_{g-1}}{\Delta_{g-1}}\\ &\qquad\qquad\quad\quad\ +\frac{\Delta_{g-1}\phi\left(r_{g-1}\right)}{2}\frac{{E_{g}}/{\Delta_{g}}-{E_{g-1}}/{\Delta_{g-1}}}{\Delta\nu_{g-1}}\Bigg].\end{split}$ (IV.12)
Here we took $R_{2,\nu_{g+1/2}}=R_{2,g}$. Notice that when $\phi=0$, Eq. IV.12 becomes the simple upwind method.
Integrating over the volume of a cell simply adds an $i$ index to the energy groups, the coefficients, and makes
$\alpha_{i}=-\sum_{j}^{\text{neighbors}}\vec{A}_{ij}\cdot\vec{v}_{ij}.$
Restructuring Eq. IV.12 results in the Doppler term for $\alpha_{i}\geq 0$:
$\begin{split}&\mathcal{D}_{g,i}[\mathbf{E}_{i}]=\\ &\Bigg[\alpha\nu_{g-1/2}\left(\frac{1-R_{2,g-1,i}}{2}\right)\left[-\frac{1}{\Delta_{g-1}}+\frac{\phi\left(r_{g-1,i}\right)}{2\Delta\nu_{g-1}}\right]\Bigg]E_{g-1,i}\\ &\Bigg[\alpha\nu_{g+1/2}\left(\frac{1-R_{2,g,i}}{2}\right)\left[\frac{1}{\Delta_{g}}-\frac{\phi\left(r_{g,i}\right)}{2\Delta\nu_{g}}\right]\\ &-\alpha\nu_{g-1/2}\left(\frac{1-R_{2,g-1,i}}{2}\right)\frac{\Delta_{g-1}\phi\left(r_{g-1,i}\right)}{2\Delta\nu_{g-1}\Delta_{g}}\Bigg]E_{g,i}+\\ &\left[\alpha\nu_{g+1/2}\left(\frac{1-R_{2,g,i}}{2}\right)\frac{\Delta_{g}\phi\left(r_{g,i}\right)}{2\Delta\nu_{g}\Delta_{g+1}}\right]E_{g+1,i},\end{split}$ (IV.13)
were $\mathbf{E}_{i}$ is the vector $\mathbf{E}_{i}\equiv(E_{1,i}\ldots E_{G,i})$, and $\mathcal{D}_{g,i}[\mathbf{E}_{i}]$ is the discrete Doppler operator.
Conversely, for $\alpha<0,\nabla\cdot v>0$, the energy groups get red-shifted. Thus, the upwind direction is towards the higher energies, and $E_{\nu_{g\pm 1/2}}$ are reconstructed by interpolation from the right side:
$\begin{split}&E_{\nu_{g}+1/2}=\frac{E_{g+1}}{\Delta_{g+1}}-\frac{\Delta_{g+1}\phi\left(r_{g+1}\right)}{2\Delta\nu_{g+1}}\left(\frac{E_{g+2}}{\Delta_{g+2}}-\frac{E_{g+1}}{\Delta_{g+1}}\right)\\ &E_{\nu_{g}-1/2}=\frac{E_{g}}{\Delta_{g}}-\frac{\Delta_{g}\phi\left(r_{g}\right)}{2\Delta\nu_{g}}\left(\frac{E_{g+1}}{\Delta_{g+1}}-\frac{E_{g}}{\Delta_{g}}\right).\end{split}$ (IV.14)
The same procedure results in the Doppler term for $\alpha_{i}<0$,
$\begin{split}&\mathcal{D}_{g,i}[\mathbf{E}_{i}]=\\ &\alpha\nu_{g-1/2}\left(\frac{1-R_{2,g-1,i}}{2}\right)\left[-\frac{1}{\Delta_{g}}+\frac{\phi\left(r_{g,i}\right)}{2\Delta\nu_{g}}\right]E_{g,i}\\ &\Bigg[\alpha\nu_{g+1/2}\left(\frac{1-R_{2,g,i}}{2}\right)\left(\frac{1}{\Delta_{g+1}}+\frac{\phi\left(r_{g+1,i}\right)}{2\Delta\nu_{g+1}}\right)\\ &-\alpha\nu_{g-1/2}\left(\frac{1-R_{2,g-1,i}}{2}\right)\frac{\Delta_{g}\phi\left(r_{g,i}\right)}{2\Delta\nu_{g}\Delta_{g+1}}\Bigg]E_{g+1,i}\\ &-\alpha\nu_{g+1/2}\left(\frac{1-R_{2,g,i}}{2}\right)\frac{\Delta_{g+1}\phi\left(r_{g+1,i}\right)}{2\Delta\nu_{g+1}\Delta_{g+2}}E_{g+2,i}.\end{split}$ (IV.15)
At the frequency boundaries $\nu_{1/2},\nu_{G+1/2}$, we impose zero frequency‑space flux, i.e., the Doppler flux across each boundary vanishes. Inserting all terms into Eq. III.2, and integrating over a single time-step yields
$\begin{split}&V_{i}E_{g,i}^{n+1}-V_{i}E_{g,i}^{n}=\\ &\Delta t\Bigl[-\left(\frac{1-R_{2,g,i}}{2}\right)E_{g,i}^{n+1}\sum_{j}^{\text{neighbors}}\vec{A}_{ij}\cdot\vec{v}_{ij}\\ &+\sum_{j}^{\text{neighbors}}D_{g,ij}\frac{\vec{A}_{ij}\cdot\left(\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right)}{\left\|\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right\|^{2}}\left(E_{g,j}^{n+1}-E_{g,i}^{n+1}\right)\\ &+cV_{i}\kappa_{P,g,i}\left(b_{g,i}U_{m,i}^{n+1}-E_{g,i}^{n+1}\right)-\mathcal{D}_{g,i}[\mathbf{E}_{i}^{n+1}]\Bigr].\end{split}$ (IV.16)
Here the superscript $n+1$ denotes values evaluated at the end of the radiation step (in the implicit scheme); quantities without a time index are evaluated at the start of the radiation step. In the Doppler term, all quantities related to the slope limiter are evaluated explicitly.
Equation IV.16 is solved together with the material internal energy density equation
$\overline{c}_{v}\frac{\partial U_{m}}{\partial t}=c\left(\sum_{g^{\prime}}\kappa_{P,g^{\prime}}E_{g^{\prime}}-\kappa_{P}U_{m}\right),$ (IV.17)
where $\overline{c}_{v}\equiv\frac{\partial T}{\partial U_{m}}\frac{\partial u_{m}}{\partial T}$ is the ratio between the heat capacity of the material and the radiation (at the material temperature). Integrating Eq. IV.17 over a single time-step results in
$\overline{c}_{v}\left(U_{m}^{n+1}-U_{m}^{n}\right)=c\Delta t\left(\sum_{g^{\prime}}\kappa_{P,g^{\prime},i}E_{g^{\prime},i}^{n+1}-\kappa_{P,i}U_{m,i}^{n+1}\right).$ (IV.18)
Rearranging Eq. IV.18 and introducing the Fleck factor (Fleck and Cummings 1971)
$f_{i}\equiv\frac{1}{1+\frac{c\Delta t}{\overline{c}_{v}}\kappa_{P,i}}$ (IV.19)
yields
$U_{m}^{n+1}=f_{i}U_{m}^{n}+\frac{1-f_{i}}{\kappa_{P,i}}\sum_{g^{\prime}}\kappa_{P,g^{\prime},i}E_{g^{\prime},i}^{n+1}.$ (IV.20)
Inserting the implicit estimation of $U_{m}$ (Eq. IV.20) into the radiation transport equation (Eq. IV.16) and moving all the implicit terms to the left side of the equality results in the following linear system:
$\begin{split}&V_{i}\left(1+c\Delta t\kappa_{P,g,i}\right)E_{g,i}^{n+1}\\ &+\Delta t\left(\frac{1-R_{2,g,i}}{2}\right)E_{g,i}^{n+1}\sum_{j}\vec{A}_{ij}\cdot\vec{v}_{ij}\\ &-\Delta t\sum_{j}D_{g,ij}\frac{\vec{A}_{ij}\cdot\left(\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right)}{\left\|\overrightarrow{CM}_{j}-\overrightarrow{CM}_{i}\right\|^{2}}\left(E_{g,j}^{n+1}-E_{g,i}^{n+1}\right)\\ &-c\Delta tV_{i}\kappa_{P,g,i}b_{g,i}\frac{1-f_{i}}{\kappa_{P,i}}\sum_{g^{\prime}}\kappa_{P,g^{\prime},i}E_{g^{\prime},i}^{n+1}+\Delta t\mathcal{D}_{g,i}[\mathbf{E}_{i}^{n+1}]\\ &=V_{i}E_{g,i}^{n}+c\Delta tV_{i}\kappa_{P,g,i}b_{g,i}f_{i}U_{m}^{n}.\end{split}$ (IV.21)
The final form of the material internal energy equation is given by inserting Eq. IV.20 into Eq. III.9, resulting in
$\frac{\partial u_{m,i}}{\partial t}=f_{i}c\left(\sum_{g^{\prime}}\kappa_{P,g^{\prime},i}E_{g^{\prime},i}^{n+1}-\kappa_{P,i}U_{m,i}^{n}\right).$ (IV.22)
Once (IV.21) is solved and the radiation field at the end of the time step is obtained, we update the internal energy via equation (IV.22). The momentum term is updated via
$\frac{{\cal M}_{i}^{n+1}-{\cal M}^{n}_{i}}{\Delta t}=-\sum_{g^{\prime}}\lambda_{g^{\prime},i}\sum_{j}^{neighbors}\vec{A}_{ij}\left(\frac{E_{g^{\prime},i}^{n+1}+E_{g^{\prime},j}^{n+1}}{2}\right)$ (IV.23)
where ${\cal M}_{i}=m_{i}\vec{v}_{i}$ is the cell’s momentum and $m_{i}$ is the cell mass. Then the kinetic energy is determined
$V_{i}K_{i}^{n+1}=\frac{\left({\cal M}_{i}^{n+1}\right)^{2}}{2m_{i}}$ (IV.24)
And finally, the total cell energy is calculated
$E_{m,i}^{n+1}=u_{m,i}^{n+1}+K_{i}^{n+1}.$ (IV.25)
IV.2 Numerical Solution of the Multigroup Equation
Equation IV.21 defines a non-symmetric sparse linear system with $N_{\rm cells}\times N_{\rm groups}$ unknowns. rich solves this linear system using an MPI-parallel, Jacobi-Preconditioned, Bi-Conjugate-Gradient-Stabilized method (BiCGSTAB), an iterative algorithm for solving non-symmetric linear systems (van der Vorst 1992). We solve the sparse linear system $\mathbf{A}\mathbf{E}=\mathbf{b}$, where $\mathbf{E}$ stacks the cell–group unknowns. Iterations terminate once all of the following hold:
  1. 1.
    Relative residual: define $\mathbf{r}\equiv\mathbf{b}-\mathbf{A}\mathbf{E}$ and require
    $\|\mathbf{r}\|\leq\varepsilon\,\|\mathbf{b}\|.$
  2. 2.
    Positivity floor: with $E_{\max}\equiv\max_{j}E_{j}$, enforce for all cells $i$ and groups $g$,
    $E_{g,i}\geq-10^{-10}\,E_{\max}.$
  3. 3.
    Component-wise residual: for every unknown $j$,
    $|r_{j}|\leq 10^{-6}\,\bigl(|A_{jj}|\,(E_{j}+10^{-5}\,E_{\max})\bigr).$
IV.2.1 Enforcing Energy Conservation
After the final linear solve, we apply a conservative defect–correction based on the component-wise residual. For an unknown $j$ representing a cell–group energy density $E_{j}$ with cell volume $V_{j}$, the balance defect in energy is $r_{j}$. We then update
$V_{j}E_{j}^{\mathrm{new}}\leftarrow V_{j}E_{j}^{\mathrm{old}}+r_{j}\quad\text{i.e.}\quad E_{j}^{\mathrm{new}}\leftarrow E_{j}^{\mathrm{old}}+\frac{r_{j}}{V_{j}}.$ (IV.26)
This post-processing step restores conservation to machine precision without additional iterations.
IV.2.2 Convergence Acceleration
In optically thick cells the radiation–material coupling becomes stiff and the block off-diagonals $c\Delta t\,\kappa_{P,g^{\prime}}$ increase the condition number, slowing BiCGSTAB. In order accelerate the convergence of the BiCGSTAB algorithm, we limit the absorption coefficients to be
$\kappa_{P,g}\leftarrow\min\left(\kappa_{P,g},\,\frac{N_{\tau}}{c\Delta t}\right)$
where we typically set $N_{\tau}=100$, which limits the local equilibration time to be $\Delta t/N_{\tau}$ while leaving the diffusive transport coefficient $\kappa_{R,g}$ unchanged. In our tests, this reduces the iteration count substantially with negligible impact on the physical solution.
V Numerical Results
V.1 Marshak Boundary Condition
The Marshak boundary condition describes the incoming flux into a medium, which is coupled to a heat bath at LTE with temperature $T_{bath}$. The boundary condition is given by (Krief and McClarren, 2024a, b; Derei et al., 2024)
$\frac{4}{c}F_{inc,g}=E_{g,b}+\frac{2}{c}F_{g,b}$ (V.1)
here $E_{g,b}$ is the group energy at the boundary, $F_{g,b}$ is the total flux on the boundary and
$F_{inc,g}=b_{g}(T_{bath})\frac{ac}{4}T_{bath}^{4}=\frac{c}{4}U_{bath,g}$ (V.2)
is the incoming group flux from the heat bath and $U_{bath,g}=b_{g}(T_{bath})aT^{4}_{bath}$. Assuming cell $i$ is a boundary cell, approximating $E_{g,b}\simeq E_{g,i}$ , inserting into Eq. V.1 and rearranging gives
$F_{g,b}=\frac{c}{2}\left(E_{g,i}-U_{bath,g}\right)$ (V.3)
In our implementation we use the implicit form of Eq. V.3. To the system described in Eq. IV.21, for a boundary cell we add $\frac{c\Delta tA_{ij}}{2}E_{g,i}^{n+1}$ , $\frac{c\Delta t}{2}A_{ij}U_{bath,g}$ to the left and right side of the equation, respectively.
V.2 Grey Diffusion
In the optically thick limit ($\lambda_{g}\to 1/3$), and taking $\kappa_{P,g}$ and $\kappa_{R,g}$ to be constant, we can recover the grey radiation-diffusion equation by summing the multigroup equations over the energy groups
$\frac{\partial E_{r}}{\partial t}+\nabla\cdot\left(\vec{v}E_{r}\right)=-\frac{E_{r}}{3}\nabla\cdot\vec{v}+\nabla\cdot\left(D\nabla E_{r}\right)+c\kappa_{P}\left(U_{m}-E_{r}\right)$ (V.4)
where $E_{r}=\sum_{g}E_{g}$. We compare the multigroup solution in this degenerate (gray) case to known analytic and semi‑analytic solutions of the grey radiation-diffusion equation. In these cases, the exact group placement is irrelevant, and only adequate coverage of the Planckian between $\nu_{1/2},\nu_{G+1/2}$ is required.
V.2.1 Non-Equilibrium Nonlinear Marshak Waves
We first test our multigroup diffusion implementation against non-equilibrium, nonlinear supersonic Marshak wave problems (Krief and McClarren (2024a); Derei et al. (2024)). These solutions are self-similar, one-dimensional and radiation-only test problems in the gray diffusion limit. We note that self-similarity methods have recently been used to extend and develop new solutions for radiation transfer (Krief (2021); Krief and McClarren (2024b); Heizler et al. (2024)) and shock hydrodynamics (Giron et al. (2021, 2023); Krief (2023)).
In each non-equilibrium Marshak wave problem, a Marshak boundary condition drives an incoming flux incoming from a heat bath with a prescribed time dependent temperature $T_{bath}\left(t\right)$, into an initially cold material $T_{m}(x,0)=0$ and a radiation field $E_{r}(x,t)\equiv 0$, with a spatially dependent density profile $\rho(x)=\rho_{0}x^{-\omega}$. The radiation temperature is
$T_{rad}=\left(\frac{E_{r}}{a}\right)^{\frac{1}{4}}.$
We consider four configurations that differ by opacity scalings and density profiles, and compare them with the analytic solutions at $t=1$ ns. Since these benchmarks are gray solutions, the opacities do not depend on frequency and are only functions of temperature and density. We note that in these benchmarks the Planck (absorption) opacity is much smaller than the Rosseland (total) opacity, so that the medium is optically thick but thermal equilibrium is not reached, while the diffusion limit is reached, so that these solutions are essentially exact transport solutions with high photon scattering and low absorption (see Krief and McClarren (2024a); Derei et al. (2024)). Regarding the numerical setup - all the cases in this section are defined on a uniform grid of 512 cells between $x\in[0,1]$ (with the exception of the last benchmark that uses 512 cells unevenly spaced), using 12 groups spanning $10^{-8}\;\text{keV}$ to $31\;\text{keV}$ logarithmically. We note that in order to test our multigroup model, we perform multigroup calculations, even though the opacities are frequency independent. The analytic similarity profiles and heat front positions are given as simple closed form analytic formulas in Krief and McClarren (2024a) and Derei et al. (2024), respectively.
Problem 1 (Test 2 from Krief and McClarren (2024a))
In this benchmark, the opacities are set to be
$\begin{split}&\kappa_{R}\left(T\right)=100\left(\frac{T}{\text{keV}}\right)^{-3}\text{cm}^{-1}\\ &\ \kappa_{P}\left(T\right)=0.001\kappa_{R}\left(T\right)\end{split}$ (V.5)
The material energy density and the density profile are uniform and given by:
$u\left(T\right)=6.860085\times 10^{14}\left(\frac{T}{\text{keV}}\right)^{4}\ \text{erg}\ \text{cm}^{-3}$ (V.6)
$\rho(x)=1\ \text{g}\ \text{cm}^{-3}$
and the driving bath temperature is
$T_{bath}\left(t\right)=1.008038\left(\frac{t}{\text{ns}}\right)^{\frac{1}{3}}\text{keV}.$ (V.7)
Figure 1: Comparison of the radiation temperature and material temperature at t=1nst=1\ $\text${ns} between the rich simulation and the analytic solution for Marshak problems 1 (left) and 2 (right). In problem 2 the absorption coefficient is large enough to enforce LTE, so $T_{\rm rad}\simeq T_{\rm mat}$.
Problem 2 (Test 3 from Krief and McClarren (2024a))
This problem is the equilbrium limit of the previous benchmark. The opacities are now set to be
$\kappa_{R}\left(T\right)=\kappa_{P}\left(T\right)=100\left(\frac{T}{\text{keV}}\right)^{-3}\text{cm}^{-1}.$ (V.8)
The material energy density and the density profile are
$u\left(T\right)=6.860085\times 10^{14}\left(\frac{T}{\text{keV}}\right)^{4}\ \text{erg}\ \text{cm}^{-3},\quad$ (V.9)
$\rho(x)=1\ \text{g}\ \text{cm}^{-3}$
and the bath temperature is
$T_{bath}\left(t\right)=1.014565\left(\frac{t}{\text{ns}}\right)^{\frac{1}{3}}\text{keV}.$ (V.10)
Figure 1 shows the material and radiation temperature as function of distance inside the slab for both benchmarks (1 & 2), and they agree nicely with the analytic solution.
Problem 3 (Test 1 from Derei et al. (2024))
In this benchmark the opacities are
$\begin{split}&\kappa_{R}\left(T,\rho\right)=40\left(\frac{T}{\text{keV}}\right)^{-1.5}\left(\frac{\rho}{\text{g}\ \text{cm}^{-3}}\right)^{1.2}\text{cm}^{-1}\\ &\ \kappa_{P}\left(T,\rho\right)=0.0025\kappa_{R}\left(T,\rho\right)\end{split}$ (V.11)
The material energy density and the density profile are spatially non-uniform and given by
$u\left(T,\rho\right)=10^{14}\left(\frac{T}{\text{keV}}\right)^{3.4}\left(\frac{\rho}{\text{g}\ \text{cm}^{-3}}\right)^{0.86}\text{erg}\ \text{cm}^{-3}\\$ (V.12)
$\rho(x)=\left(\frac{x}{\text{cm}}\right)^{\frac{20}{19}}\ \text{g}\ \text{cm}^{-3}$
and the bath temperature is
$T_{bath}\left(t\right)=1.0470478\left(\frac{t}{\text{ns}}\right)^{\frac{86}{57}}\text{keV}$ (V.13)
Problem 4 (Test 3 from Derei et al. (2024))
The opacities are now set to be
$\begin{split}&\kappa_{R}\left(T,\rho\right)=2\left(\frac{T}{\text{keV}}\right)^{-4.5}\left(\frac{\rho}{\text{g}\ \text{cm}^{-3}}\right)^{1.9}\text{cm}^{-1}\\ &\kappa_{P}\left(T,\rho\right)=0.05\kappa_{R}\left(T,\rho\right)\end{split}$ (V.14)
The material energy density and the density profile are (notice that the density profile diverges at $x=0$)
$u\left(T,\rho\right)=10^{14}\left(\frac{T}{\text{keV}}\right)^{6}\left(\frac{\rho}{\text{g}\ \text{cm}^{-3}}\right)^{0.7}\ \text{erg}\ \text{cm}^{-3}$ (V.15)
$\rho(x)=\left(\frac{x}{\text{cm}}\right)^{-\frac{40}{139}}\ \text{g}\ \text{cm}^{-3}$
and the bath temperature is
$T_{bath}\left(t\right)=1.01008116\left(\frac{t}{\text{ns}}\right)^{\frac{14}{139}}\text{keV}.$ (V.16)
Unlike the previous benchmarks, the divergence of the density at the origin required us to use a refined grid in which the Voronoi mesh generating points are located at
$x_{i}=10^{-5}+2.24\times 10^{-2}\left(1.0075^{i}-1\right)\ \text{cm}.$ (V.17)
Figure 2 shows the material and radiation temperature as function of distance inside the slab for both benchmarks (3 & 4), and they agree nicely with the analytic solution.
Figure 2: Comparison of the radiation temperature and material temperature at t=1nst=1\ $\text${ns} between the rich simulation and the analytic solution for Marshak problems 3 (left) and 4 (right).
V.2.2 Mach 2 Radiative Shock
In order to test our coupling of the radiation transport to the hydrodynamics, we run the Mach 2 test problem (Ferguson et al., 2017; Lowrie and Edwards, 2008). This is a one-dimensional, non-equilibrium, radiation-hydrodynamics test problem in the grey diffusion approximation.
The problem consists of two semi-infinite regions of a perfect gas with the same adiabatic index $\gamma$. The problem is defined using the values of the gas at -$\infty$ for the left region $\rho_{\ell},T_{\ell},v_{\ell}$ and at $\infty$ for the right region $\rho_{r},T_{r},v_{r}$.
A shock front develops between the two regions, and in the shock frame the steady state solution has a semi-analytic solution.
The initial conditions for the left region $x<0$ are:
$\rho_{\ell}=5.45887\times 10^{-13}\ \text{g}\ \text{cm}^{-3}$ (V.18)
$T_{\ell}=100\ \text{K},\ v_{\ell}=2.3547\times 10^{5}\ \text{cm}\ \text{s}^{-1}.$
And for the right region $x>0$
$\rho_{r}=1.2479\times 10^{-12}\text{g}\ \text{cm}^{-3}$ (V.19)
$T_{r}=207.757\ \text{K},\ v_{r}=1.03\times 10^{5}\ \text{cm}\ \text{s}^{-1}.$
The properties of the gas are given by
$\begin{split}&\gamma=\frac{5}{3},\ c_{v}=1.2472\times 10^{8}\ \text{erg}\ \text{K}^{-1}\ \text{g}^{-1}\\ &\kappa_{R}=0.848902\ \text{cm}^{-1},\ \kappa_{P}=3.93\times 10^{-5}\ \text{cm}^{-1}\end{split}$ (V.20)
The simulation is performed using a uniform grid of 256 cells between x=1000x=-1000 cm and $x=2000$ cm, with $5$ energy groups between $10^{-3}\text{eV}$ and $1\text{eV}$ spaced logarithmically. We evolve the system until $t=0.02$ seconds, by then the system has evolved to the steady state solution. Figure 3 shows the gas temperature and radiation energy density profiles along with the semi‑analytic solution.
Figure 3: The gas temperature (left) and radiation energy density (right) for the Mach 2 benchmark along with the semi-analytic solution.
V.3 Multigroup benchmarks
V.3.1 Densmore Multigroup tests
We validate our multigroup diffusion module by comparing it with the benchmarks presented in Densmore et al. (2012). The benchmarks consists of one-dimensional, radiation-only, non-linear multigroup problems. The reference data from Densmore et al. (2012) (see also Steinberg and Heizler (2022b, 2023)) were digitized from their figures, which were obtained using a Monte Carlo method that remains reliable in low-opacity regimes.
All of the setups consist of a Marshak boundary condition at the left boundary located at $x=0$, with a bath temperature of $1\;\text{keV}$, an initially cold ideal gas in the domain $x\in[0,L]$, and an open boundary at $x=L$. The parameters of the equation of state are:
$\gamma=1.4,\quad c_{v}=\frac{10^{15}}{\text{keV}}\ \text{erg}\ \text{cm}^{-3}.$ (V.21)
The opacity has the following spatial, temperature and frequency dependence:
$\kappa_{R,g}\left(x,T\right)=\kappa_{P,g}(x,T)=\frac{\kappa_{0}\left(x\right)}{\nu_{g}^{3}\sqrt{k_{B}T}}\text{cm}^{-1}$ (V.22)
and the different benchmarks differ in their corresponding $\kappa_{0}(x)$. All simulations are run with 32 logarithmically spaced groups between $10^{-4}\;\text{keV}$ and $100\;\text{keV}$ with a uniform grid of 128 cells between $[0,L]$.
In the first three runs, $\kappa_{0}(x)$ was constant and equal to
$\kappa_{0}=10\;\text{keV}^{3.5},1000\;\text{keV}^{3.5},100\;\text{keV}^{3.5}$ (V.23)
The system length $L$ differs between the runs and is respectively
L=5, 3, 1cmL=5,\;3,\;1\ $\text${cm} (V.24)
Figure 4: Comparison of the gas temperature at $t=1$ ns between rich and the data taken from Densmore et al. (2012) for the cases $\kappa_{0}=10\;\text{keV}^{3.5}$ (left) and $\kappa_{0}=100\;\text{keV}^{3.5}$ (right)
For the fourth run $\kappa_{0}(x)$ was a step function, jumping from an optically thin region to an optically thick region
$\kappa_{0}(x)=\begin{cases}10\;\text{keV}^{3.5}&,x<2\\ 1000\;\text{keV}^{3.5}&,x\geq 2\end{cases}$ (V.25)
with a system of $L=3$ cm.
Figures 4 and 5 show the gas temperature at the time $t=1\;\text{ns}$ for the different benchmarks. Overall, even though the benchmarks are compared to a Monte-Carlo method, the agreement is good for all of the tests.
Figure 5: Comparison of the gas temperature at $t=1$ ns between rich and the data taken from Densmore et al. (2012) for the medium opacity case $\kappa_{0}=1000\;\text{keV}^{3.5}$ (left) and the step function opacity as presented in Eq. V.25 (right).
V.3.2 Doppler Term
To validate the Doppler term, we construct a zero-dimensional test problem. Starting from Eq. II.1, we consider the optically thick limit and omit advection, diffusion, and emission–absorption terms, isolating the Doppler contribution:
$\frac{\partial E_{\nu}}{\partial t}=-\left(\frac{1}{3}E_{\nu}-\frac{1}{3}\frac{\partial\left(\nu E_{\nu}\right)}{\partial\nu}\right)\nabla\cdot\vec{v}$ (V.26)
This configuration constitutes an artificial variant of the radiation–diffusion equation. We include the hydrodynamic dependence through $\nabla\cdot\vec{v}$, but intentionally omit it from the advection term in order to isolate the effect of the Doppler contribution.
In order to simplify the notation, we denote $K\equiv-\frac{\nabla\cdot\vec{v}}{3}$ and rewrite Eq. (V.26) as:
$\frac{1}{K}\frac{\partial E_{\nu}}{\partial t}+\frac{\partial\left(\nu E_{\nu}\right)}{\partial\nu}=E_{\nu}.$ (V.27)
This can be simplified to be:
$\frac{1}{K}\frac{\partial E_{\nu}}{\partial t}+\nu\frac{\partial E_{\nu}}{\partial\nu}=0.$ (V.28)
Assuming $K$ is constant, equation (V.28) is known to have a solution in the form (Polyanin et al., 2001)
$E\left(\nu,t\right)=\Phi\left(\nu e^{-Kt}\right),$ (V.29)
where $\Phi$ is a function that is given by the initial conditions. By equating the two sides at time $t=0$
$E(\nu,0)=\Phi(\nu)$ (V.30)
we can recover the general solution as a function of frequency and time
$E\left(\nu,t\right)=E\left(\nu e^{-Kt},0\right).$ (V.31)
For this benchmark, we set $D_{g}=0$, $\kappa_{P,g}=0$ and $\lambda=\frac{1}{3}$ in rich, and run only the radiation transport (no hydrodynamic evolution).
The test consists of two identical box-shaped cells. The left cell is stationary with $\vec{v}_{\ell}=\vec{0}$, while the right cell is assigned a velocity of $\vec{v}_{r}=(10^{9},0,0)\;\text{cm}\cdot\text{s}^{-1}$. This configuration generates a non-zero $\nabla\cdot\vec{v}$, with opposite signs in the two cells, thereby creating a controlled expansion in the left cell and compression in the right cell, ideal for testing the Doppler term.
The initial radiation profile (which is arbitrary), is given by a truncated Planck function with a temperature of 1 keV
$E\left(\nu,0\right)=\begin{cases}B\left(\nu,T=1\text{keV}\right)&1.12\;\text{keV}<\nu<8.12\;\text{keV}\\ 0&\text{else.}\end{cases}$ (V.32)
To assess the impact of slope limiting (see Eqs. IV.13 and IV.15), we compare a first-order upwind scheme with $\phi=0$ to the Superbee limiter (Eq. IV.10). As illustrated in Fig. 6, the Superbee limiter yields markedly lower numerical diffusion than the upwind approach in both compressive and expansive regimes, resulting in a more accurate solution of the Doppler term.
Figure 6: Comparison between the rich and the analytic solution for the case of the upwind method (left) and the Superbee slope-limiter (right) at a time $t=40$ ns. Using the Superbee limiter drastically reduces the numerical diffusion compared to the simple upwind method in both the compression and expansion cases.
V.3.3 Convergence Acceleration
The improved convergence method described in section IV.2.2 is evaluated using an astrophysically motivated benchmark. Specifically, we model a Marshak wave under physical conditions characteristic of Tidal Disruption Event simulations (Steinberg and Stone, 2024). The setup consists of a 1D domain with a Marshak boundary condition imposed at one end, where a bath temperature of $T_{bath}=10^{5}\;\text{K}$ drives radiation into cold gas of density $\rho=10^{-7}\;\text{g/cc}$. The gas follows an ideal-gas equation of state with adiabatic index $\gamma=5/3$ and a heat capacity of $C_{v}=10^{8}\rho\;\text{erg}\;\text{cm}^{-3}\;\text{K}^{-1}$. For the opacity we assume a free-free opacity described by $\kappa_{P,\nu}=\kappa_{R,\nu}=3.7\cdot 10^{8}n^{2}\nu^{-3}T^{-1/2}\cdot(1-\exp{(-h\nu/k_{B}T)})\;\text{cm}^{-1}$, assuming the gas is completely ionized at all times and all of the quantities are in CGS units. We evolve the system until a time of $t=10^{3}\;\text{s}$ on a uniform grid $x\in(0,\;10^{11})\;\text{cm}$ divided uniformly into 256 cells once with our convergence acceleration using $N_{\tau}=100$ and once without it. Figure 7 presents the gas temperature and the radiation temperatures in two energy groups as functions of position for both simulations. The profiles are nearly identical, confirming that the convergence-acceleration method does not alter the physical solution. However, the computational cost differs substantially: the run without acceleration required 923 s, whereas the accelerated run completed in only 82 s.
Figure 7: Temperature profiles of the gas (left) and of the radiation (right) in two energy groups (14 eV and 94 eV) for simulations conducted with and without the convergence-acceleration scheme. The overlapping curves confirm that the method maintains solution accuracy. The accelerated run reduces the computational cost by more than an order of magnitude, completing in 82 s compared to 923 s.
V.4 Tidal Disruption Event
To demonstrate the astrophysical applicability of the new multigroup radiation–diffusion module, we employ it in a full, self-consistent tidal disruption event (TDE) simulation. We simulate the parabolic approach of a star with mass $M_{\star}=0.5M_{\odot}$ and radius $R_{\star}=0.47R_{\odot}$ towards a $M_{\bullet}=10^{4}M_{\odot}$ black hole. The star is quickly torn apart by the black hole’s tidal field, and its deconfined debris flies out to radii $\sim R_{\star}(M_{\bullet}/M_{\star})^{2/3}$ before the most bound material turns around at its apocenter. As stellar debris returns to pericenter, it begins dissipating orbital energy into heat through shocks. Eventually, a subset of this dissipated energy is radiated, producing the luminous nuclear flares seen today by time-domain observers (van Velzen et al., 2020). The details of shock dissipation, debris circularization, and radiative emission all remain hotly debated at present (Hayasaki et al., 2013; Piran et al., 2015; Shiokawa et al., 2015; Bonnerot and Stone, 2021; Bonnerot et al., 2021; Andalman et al., 2022; Steinberg and Stone, 2024; Huang et al., 2024; Andalman et al., 2025), and this proof-of-principle simulation does not aim at resolving major open questions in TDE physics. Because the simulation has the exact astrophysical parameters as the grey-rich simulations of (Martire et al., 2025), we generally expect a similar result: much weaker dissipation and slower circularization than what is seen in supermassive black hole TDEs; emission powered by dissipation at the nozzle shock that is reprocessed through a radiation-driven wind; near-Eddington peak luminosities. The main advantage of this work relative to past end-to-end TDE simulations is the use of multigroup (rather than grey) radiation transfer, which allows us to self-consistently predict light curve evolution in different frequency bands.
For this simulation, the maximum resolution reached was $18\times 10^{6}$ cells, divided across 256 cores. The gas opacities are interpolated from tabular data calculated using the STAR atomic code, which employs the super-transition-array method (Krief et al., 2018a; Krief and Feigel, 2015a, b; Krief et al., 2018b). For the radiation, we used 10 energy groups, presented in Table 1. Together these groups span the full range of energies in which observed TDEs have been seen to emit quasi-thermally [^1] I.e. we ignore the radio and gamma-ray photons that are produced via non-thermal processes related to outflows in a subset of TDEs..
Name $\nu_{low}\ [eV]$ $\nu_{high}\ [eV]$
i 0.01 1.6767
r 1.6767 2.4796
b 2.4796 3.3508
U 3.3508 4.1327
UVW2 4.1327 7.7488
H 7.7488 15
He 15 100
Super Soft (SS) 100 300
Soft X-ray 300 1000
Hard X-ray 1000 20000
Table 1: Photon energy groups used in the multigroup TDE simulation. Most thermal emission from TDEs has been observed in our r-, b-, U-, UVW2, $S$- and Soft X-ray bands. Theoretical and observational arguments suggest that the majority of TDE emission may lie in unobservable EUV wavelengths, parametrized here as “H-” and “He-” bands.
A more detailed description of the setup, initial conditions, and adaptive mesh refinement criteria is given in Martire et al. (2025). The simulation is run until it reaches a time of 2.1 fallback times ($t_{\rm fb}=2.57\ \text{days}$), at which point it has achieved its peak bolometric luminosity.
Figure 8 shows a density slice through the orbital plane of the simulation at $t=0.42\ t_{\rm fb}$. At this point, the most tightly bound debris has just begun to return to pericenter, and dissipation in the nozzle shock is beginning to produce the rising phase of the light curve.
Figure 8: Mass density slice in the orbital plane at $t=0.42\ t_{\rm fb}\approx 1\ \text{day}$ after the star was disrupted. The stellar debris has begun falling back, forming an accretion flow around the black hole. Gas densities are labeled according to the color bar in the figure, and are highest in the debris stream.
We extract the luminosity in each energy group along a given viewing angle $\vec{\Omega}$ (the angular grid is generated using HEALPix with $N_{\rm side}=8$; Górski et al. 2005)
$L_{g}(r_{\vec{\Omega}},\vec{\Omega},t)=4\pi r_{\vec{\Omega}}^{2}cE_{g}\left(r_{\vec{\Omega}},\vec{\Omega},t\right)$ (V.33)
by identifying the distance $r_{\vec{\Omega}}$ along a ray originating at the black hole, where the outward group optical depth reaches
$\tau_{\rm g}=\intop_{r_{\Omega}}^{\infty}\kappa_{R,g}dr=\frac{2}{3}.$ (V.34)
Figure 9: Mollweide sky maps of the Optical+UV luminosity (left) and the Soft+Hard X-ray band luminosities (right), at $t=0.4\ t_{fb}\approx 1\ \text{day}$. At this early stage of evolution, very little luminosity is being produced, but hints of later trends have already emerged: optical+UV emission is primarily emitted in the orbital plane where the reprocessing column density is highest, while X-ray emission is suppressed in the orbital plane for the same reason.
Figure 10: Mollweide sky maps of the Optical+UV luminosity (left) and the Soft+Hard X-ray band luminosities (right), at $t=0.78\ t_{\rm fb}\approx 2\ \text{days}$, near the peak of the early X-ray flash. X-ray luminosities have reached near-maximum values, but are suppressed within the orbital plane where they are preferentially absorbed by stellar debris. Conversely, optical/UV emission is overwhelmingly concentrated in these same planar lines of sight.
Figure 11: Mollweide sky maps of the Optical+UV luminosity (left) and the Soft+Hard X-ray band luminosities (right), at $t=2.08\ t_{\rm fb}\approx 5.3\ \text{days}$. At this late stage, an optically thick photosphere has emerged across $4\pi$ Sr, heavily absorbing X-rays along all lines of sight and producing a more (though not fully) isotropic angular distribution of optical/UV emission.
Figure 12: Mollweide sky maps of the ratio between Optical+UV to X-ray luminosity (left) and Optical+UV to our Helium band luminosity (right), at $t=0.79\ t_{fb}\approx 2\ \text{day}$, near peak X-ray luminosity. The optical and UV luminosity is strongest along the orbital plane, where dense material efficiently reprocesses the radiation. In contrast, the X-ray emission originates closer to the black hole and escapes preferentially in directions away from the plane, where the column density is lower and attenuation is reduced.
We further define the angle-averaged group luminosity to be
$L_{\rm g}(t)\equiv\frac{1}{4\pi}\oint_{4\pi}L_{g}(r_{\vec{\Omega}},\vec{\Omega},t)d\Omega.$ (V.35)
In Figs. 911 we show isotropic equivalent Mollweide skymaps of band luminosities at t=0.40, 0.79,t=0.40,\,0.79, and $2.08\,t_{\rm fb}$, respectively. Figure 12 also presents ratio maps (Optical+UV/X-ray and Optical+UV/He) at $t=0.79\,t_{\rm fb}$. Near peak luminosity, the Optical+UV emission is strongest along the disrupted star’s orbital plane, consistent with reprocessing of an ionizing continuum in dense debris, whereas the X-ray bands peak at higher inclinations off the plane, where the reprocessing column density is lower (see especially Figs. 1012). At $t=2.08\,t_{\rm fb}$ a localized X-ray bright spot appears (Fig. 11), density and temperature slices indicate it arises from hot gas breaking out of the inner disk (Figs. 1415).
We further quantify the observable features of this simulated TDE by plotting sky-averaged multiband light curves, in Fig. 13. The sky-averaged light curves exhibit a clear X‑ray flash, peaking at $\approx 2.5$ days after disruption (Fig. 13). This result appears consistent with earlier predictions of similar X-ray flashes in supermassive black hole TDEs from Steinberg and Stone (2024), which appear to have been observed in the (supermassive) TDE AT 2022dsb (Malyali et al., 2024). In contrast to Steinberg and Stone (2024), we have made self-consistent predictions for the time-dependent X-ray luminosity in this paper using our novel multigroup FLD transfer module; with grey simulations, spectral predictions can only be made in post-processing, which may not be self-consistent in situations like TDE circularization where (frequency-dependent) radiation pressure dominates over purely hydrodynamical forces.
Both this multigroup simulation and the post-processing of grey RHD results in Steinberg and Stone (2024) suggest that early X-ray flashes can be produced by the activity of the nozzle shock at times (and viewing angles) when the column density of colder absorbing debris is not yet too large. This X-ray flash merits further investigation: if detected in future samples of TDEs, it could serve as a “starting gun,” timing the onset of mass return to pericenter. As many of the thorniest theoretical questions concerning TDEs are downstream from long-standing uncertainties concerning the duration of the circularization process, a concrete measurement of the time when circularization begins would be highly valuable.
In agreement with some prior theoretical work (Lu and Kumar, 2018; Dai et al., 2018; Wen et al., 2020), the bulk of our flare’s bolometric luminosity emerges at unobservable EUV wavelengths (i.e. our H and He frequency groups, to which the Milky Way’s ISM is optically thick in absorption). These frequency groups track the X-ray band evolution at early times (up to $\approx 1t_{\rm fb}$, and peak at timescales comparable to the early X-ray peak, but they do not suffer the same rapid, multiple order of magnitude drop seen in X-rays, likely due to intermediate temperature gas (i.e. gas that thermally emits in the EUV) at the color surface. At even longer wavelengths, Fig. 13 shows a monotonic increase in optical and near-UV luminosity as a function of time.
Figure 13: Light curves for the different energy groups. Most radiation emerges in UV wavelengths; primarily in the unobservable EUV bands (H and He). UV bands (including Swift UVW2) reach peaks or plateaus near $t\approx 1t_{\rm fb}$, while optical wavelengths monotonically rise for the duration of the simulation. The earliest X-ray flash (at $t\approx 0.6~{\rm day}$) is likely an unphysical artifact of under-resolving the tip of the stream as it comes back to pericenter (Martire et al., 2025), but the broader X-ray flare from $\approx 1$ to $\approx 2.5$ days is more likely physical, and agrees qualitatively with both the simulation of Steinberg and Stone (2024) and observations of the TDE AT 2022dsb (Malyali et al., 2024). This paper’s multigroup simulations are the first self-consistent prediction of an early X-ray flash in an end-to-end simulation, suggesting that these flares may offer observational insight into both circularization mechanisms and viewing angles.
The dissipation dynamics in this simulation are complex, which we illustrate in Fig. 14 using midplane density and temperature slices taken towards the end of the simulation. Visual inspection of the density map shows that even though little circularization has occurred (in agreement with the grey RHD IMBH simulation of Martire et al. 2025) due to the general weakness of the nozzle shock (Guillochon et al., 2014; Fitz Hu et al., 2025; Andalman et al., 2025), the flow has become turbulent. Gas temperatures are highest in the central regions of the eccentric accretion flow, although a thin band or stream of high-temperature gas just upwards (in-plane, along the positive $y$-axis), as can be better seen in the Fig. 15 zoom-in. This hot stream is ultimately responsible for the limited remaining late-time X-ray emission, as can be seen by comparing to the Mollweide plot in Fig. 11. Early-time emission from TDEs likely exhibits a complex viewing angle dependence, especially for harder photon energies most susceptible to absorption and reprocessing.
A deeper investigation of multigroup simulations of TDEs will be presented in Giron et al in prep, but the pilot simulation shown here illustrates the potential of our new module for making directly testable predictions about TDE emission. While this paper was in final stages of preparation, Huang et al. 2025b published a different multigroup RHD simulation of a TDE. While these two papers are substantially different from each other (in contrast to our end-to-end IMBH TDE simulation, Huang et al. 2025b present a SMBH TDE simulation that skips over stream dynamics by using a mass injection scheme similar to that in Huang et al. 2025a), the conclusions about the angular dependence of different frequency groups appear qualitatively similar. Notably, both works confirm the earlier prediction of Steinberg and Stone (2024) regarding early X-ray emission originating from shock dissipation.
(a) a
(b) b
Figure 14: (a) Density slice in the orbital plane at $t=2.08t_{\rm fb}(\approx 5.3\ \text{day})$ after the star was disrupted. (b) Slice in the orbital plane of the X-ray energy density, taken at the same time. The pericentric nozzle shock has generated a complex and turbulent flow, even though little circularization has occurred in absolute terms. A thin, high-temperature band parallels a density jump extending upwards (along the in-plane $y$-axis) from the nozzle shock.
Figure 15: Gas temperature slice in the orbital plane at $t=2.08\ t_{fb}(\approx 5.3\ \text{days})$, zoomed on the pericenter and “breakout” region. The stream of hot gas visible on the positive $y$-axis is the origin of the residual late-time X-ray emission visible in Fig. 11, illustrating the complex viewing angle dependence of high-energy radiation at early times in TDEs.
VI Conclusions
We have modified the moving-mesh radiation hydrodynamics code rich to incorporate a multigroup module for radiation transfer. This builds on our previous work, adding grey mixed frame FLD radiation transfer into rich, the results of which were previously presented in Steinberg and Stone (2024). The updated version of rich is publicly available, in keeping with our longstanding policy of accessible code development. rich is now the only moving-mesh RHD code capable of transporting radiation using multigroup photons, creating a unique capability for astrophysical problems where (i) radiation is dynamically important and (ii) the relevant radiation field spans orders of magnitude in photon energy. The updated rich will also be valuable for general problems in astrophysical hydrodynamics where broadband spectral predictions are useful, reducing the need for post-processing of hydrodynamic simulations with dedicated multigroup radiation transfer codes.
Our implementation of multigroup FLD into rich expands the capabilities of moving-mesh RHD at the cost of greater complexity and computational expense. The multi-step RHD algorithm is described in more detail in §IV, but in brief, the code first updates the hydrodynamic variables, then moves the Voronoi mesh to follow the bulk fluid flow, and then updates the FLD radiation equation [^2] This overview is slightly oversimplified; for example, radiation advection occurs during the hydrodynamics sub-step.. Finally, radiation forces update the gas momentum and energy variables. To mitigate the greater computational cost of performing the RHD in a multigroup way, we have implemented a novel scheme to accelerate convergence in optically thick cells, with minimal loss of accuracy (as can be seen in Fig. 7).
We have validated both the grey and the multigroup RHD modules in rich in a series of test problems, presented in §V.2 and §V.3, respectively. In general, the numerical solutions from rich achieve excellent agreement with previously known self-similar test problems and analytic jump conditions for radiative shocks, as well as a zero-dimensional test problem we have constructed to probe the Doppler term in multigroup RHD.
As a final test for multigroup rich, we have gone beyond test problems with analytic solutions and have run the first end-to-end simulation of a TDE in multigroup RHD. Our simulation follows the disruption of a lower main sequence star by a $M=10^{4}M_{\odot}$ IMBH, using astrophysical parameters identical to the grey RHD simulations recently presented in Martire et al. (2025). This simulation produced multi-band light curves across 10 different frequency groups, and runs for a little over 2 fallback times. Although we will defer more detailed explorations of multigroup TDE simulations for future work (Giron et al. in prep), this proof-of-principle simulation is already quite interesting. The bulk hydrodynamics and bolometric light curve appear broadly consistent with the slow circularization seen in Martire et al. (2025), but the emergent radiation field exhibits a complex angular dependence and frequency evolution. Optical and near-UV bands rise monotonically but never make up more than $\sim 1\%$ of the bolometric emission, which is released primarily in the unobservable EUV (although we note that the far-UV Swift UVW2 band does eventually reach $\sim 10\%$ of the bolometric luminosity).
The light curve evolution at short wavelengths is richer: both soft and hard X-ray bands rise rapidly in tandem with bolometric luminosity, and at the peak of the bolometric light curve (achieved around $t\approx 1t_{\rm fb}$), a few tens of percent of the emitted radiation is in X-rays. However, X-rays can only escape along lines of sight with low gas column density, and the partial isotropization of the stellar debris quickly closes the window on X-ray emission. The X-ray light curve peaks near $1t_{\rm fb}$ and then enters a rapid, multiple orders of magnitude decline, although we anticipate it will eventually rise again due to the onset of central accretion (not treated in this simulation). This early time, pre-peak X-ray flash qualitatively agrees with the SMBH TDE predictions of Steinberg and Stone (2024), although here we have produced our spectral predictions in a more self-consistent way. This early X-ray flash, which ultimately originates from shock dissipation in the first stages of the circularization process, may already have been seen in one TDE [^3] Immediately prior to completion of this paper, an analysis of the eROSITA X-ray selected TDE sample appears to have uncovered a second TDE with an X-ray flash prior to optical peak, SRGe J131014.7+444319 (Zhang et al., 2025). (Malyali et al., 2024), and could provide direct observational constraints on the circularization efficiency if it could be studied in more detail in a broader TDE sample.
We note two limitations of the present comoving-frame FLD formulation. Because a conserved total (gas+radiation) energy is defined only in an inertial frame, the comoving-frame equations do not admit a manifestly conservative form for the combined energy budget. This is especially undesirable for dynamically evolving meshes such as the moving Voronoi grid used by rich (cf. Krumholz et al. 2007a). In addition, we truncate the system at first order in $v/c$, neglecting $\mathcal{O}(v^{2}/c^{2})$ (and higher) corrections that can become leading-order in the dynamic-diffusion regime and in nonuniform, time-dependent flows (Krumholz et al., 2007a; He et al., 2024, § 3.1).
In the future, we hope that a publicly available, multigroup moving-mesh RHD code may provide broad utility in astrophysics. We anticipate that the radiation transfer module in rich will be particularly useful for simulating astrophysical problems where (i) radiation forces become dynamically important and (ii) some combination of large dynamic range and highly supersonic flow pose severe challenges for more standard numerical hydrodynamics schemes. Multigroup rich will be especially applicable for the subset of these problems where broad-band light curves or spectral predictions are desired. Within our own collaboration, we anticipate extending the pilot multigroup TDE simulation presented here to a broader range of TDE parameters. In the very near future, both the Vera Rubin Observatory (Bricman and Gomboc, 2020) and the Israeli-NASA-DESY ULTRASAT UV survey (Shvartzvald et al., 2024) will likely discover thousands of TDEs, but the analysis of these flares could be gravely limited by the lack of first-principles models. A broader parameter survey of multigroup RHD TDE simulations may be a necessary step in realizing the scientific potential of TDEs in this decade.
Acknowledgement
We thank P. Chang, P. Martire, and E.M. Rossi for fruitful discussions. NCS acknowledges support from the Binational Science Foundation (grant No. 2020397) and the Israel Science Foundation (Individual Research Grant No. 2414/23). All the simulations were performed on the ICPL cluster at the Racah Institute of Physics.
Appendix A Computing Group Planck Integrals
The normalized group Planck integrals $b_{g}$ were computed using a truncated polylogarithmic and Taylor series taken from Equations 32, 38 in Clark (1987) , respectively. Taking the first 5 exponentials in the polylogarithmic series, the first 9 terms in the Taylor series and stitching them together at $x=2$ results in an accuracy of at least $10^{-5}$ (see Fig. 3 in Clark (1987)). Concretely, we calculate $b_{g}$ using the following expression.
$b_{g}(T)=\Psi(\nu_{g+1/2}/k_{B}T,\ \nu_{g-1/2}/k_{B}T)$ (A1)
where
$\Psi\left(x_{1},x_{2}\right)=\begin{cases}1+\frac{15}{\pi^{4}}\left(\Pi\left(x_{2}\right)-\Pi\left(x_{1}\right)\right),&x_{2}>2>x_{1}\\ \frac{15}{\pi^{4}}\left(\Pi\left(x_{2}\right)-\Pi\left(x_{1}\right)\right),&\text{else}\end{cases}$ (A2)
$\Pi\left(x\right)=\begin{cases}-x^{3}\sum_{l=1}^{5}\frac{e^{-lx}}{l}-3x^{2}\sum_{l=1}^{5}\frac{e^{-lx}}{l^{2}}-6x\sum_{l=1}^{5}\frac{e^{-lx}}{l^{3}}-6\sum_{l=1}^{5}\frac{e^{-lx}}{l^{4}},&x>2\\ \frac{x^{3}}{3}-\frac{x^{4}}{8}+\frac{x^{5}}{60}-\frac{x^{7}}{5040}+\frac{x^{9}}{272160},&\text{else}\end{cases}$ (A3)
A detailed survey and analysis of the evaluation of incomplete Planck integrals was recently published by Lewis and McClarren (2025).